14. The NPV and payback period What information does the payback period provide? Suppose you are evaluating a project with the expected future cash inflows shown in the following table. Your boss has asked you to calculate the projects net present value (NPV). You dont know the projects initial cost, but you do know the projects regular, or conventional, payback period is 2.50 years If the projects ~WACCs is 10%, the projects NPV (rounded to the nearest dollar) is: Year Cash Flow Year $325,000 Year 2 $425,000 Year 3 $400,000 Year4 $425,000 O $287,501 O $345,001 O $230,001 o $273,126 Which of the following statements indicate a disadvantage of using the regular payback period (not the discounted payback period) for capital budgeting decisions? Check all that apply The payback period does not take the projects entire life into account. The payback period does not take the time value of money into account. The payback period is calculated using net income instead of cash flows
Answer #1

Payback period=Last period with a negative cumulative cash flow+(Absolute value of cumulative cash flows at that period/Cash flow after that period).

Hence intial investment=325000+425000+(0.5*400,000)=$950,000

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=325000/1.1+425000/1.1^2+400,000/1.1^3+425000/1.1^4

=$1237500.85

NPV=Present value of inflows-Present value of outflows

=$1237500.85-950,000
=$287501(Approx)

Payback period unlike the discounted payback does not take into account the time value of money unlike the discounted payback method.It only takes into account the period within which the initial investment is recovered.

Hence the correct option is A and B.